Obtain a family of solutions for the homogeneous differential equation
1
Expert's answer
2021-09-12T23:52:49-0400

a)


"y''-8y'+12y=0"

Auxiliary (characteristic) equation


"r^2-8r+12=0"

"(r-2)(r-6)=0"

"r_1=2, r_2=6"

The family of solutions of the given differemtial equation is


"y=C_1e^{2t}+C_2e^{6t}"

b)


"y''+10y'+25y=0"

Auxiliary (characteristic) equation


"r^2+10r+25=0"

"(r+5)^2=0"

"r_1=r_2=-5"

The family of solutions of the given differemtial equation is


"y=C_1e^{-5t}+C_2te^{-5t}"

c)


"y''-8y'+65y=0"

Auxiliary (characteristic) equation


"r^2-8r+65=0"

"D=(-8)^2-4(1)(65)=-196<0"

"r_{1,2}=\dfrac{8\pm\sqrt{-196}}{2(1)}=4\pm7i"

The family of solutions of the given differemtial equation is


"y=C_1e^{4t}\cos(7t)+C_2e^{4t}\sin(7t)"